OpenStudy (anonymous):

Can someone help with this equation with a calculator: 3x+19=32 (round to the nearest hundredth)

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x = 13 }\) \( \color{Black}{\Rightarrow x = \Large{13 \over x} }\) You should use fractions when in need ;)

Parth (parthkohli):

Oops

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = \Large{13 \over 3} }\)

OpenStudy (anonymous):

how do you do it like this: 3*___ +19= 32. easily

OpenStudy (anonymous):

Think.

OpenStudy (anonymous):

The answer is 4.3333333333333333333333333333

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

\[3x+19=32 subtract 19 from 32 and you get 13 divde 13 by 3\]

OpenStudy (anonymous):

ok so how about 7x-8=17

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3.57 is the answer

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

\[7x - 8 = 17 you add 8 \to 17 and you get 25 then divide \it by 7\]

OpenStudy (anonymous):

k cool this might be the last one but... -2+4x=-23

OpenStudy (anonymous):

i got -23.5

OpenStudy (anonymous):

Its 5.25

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

\[You add 2 \to 23 becasue they are negative they become positive and you 25 divide by \]

OpenStudy (anonymous):

ohhhhhhhhh ok is it ok i keep doing this until i get it?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8x-3=-64

OpenStudy (anonymous):

i got -24

OpenStudy (anonymous):

Its 8.375 you add 3 to 64 and divide it by 8

OpenStudy (anonymous):

ohh ok now 27=58-6x

OpenStudy (anonymous):

i got 13.83

OpenStudy (anonymous):

is that right?